if x = the the radical sign 3, y=2 and z = 1/2, then x square - 5yz+yto the 2= 1. 1 B. 2 C. 4 D. 7 B is the answer.
If x 3^, y=2, and z =1/2, the x^- 5yz +y^ = ?
^I hope this means the square sign.

Answer by Boreal(15235)

x^2=3
-5yz=-5*2*1/2=-5
and y^2=4
3-5+4=2 the answer. B
I'm not clear on the second.
^2 is square sign
^3 is cube
^ means exponent is being named.